June earns $3 an hour for the first 40 hours, then $4.50 an hour for overtime. How many hours must she work in order to make at least $165?
Answer by Mathpassionate(25)

Let's use:
E = number of extra(overtime)hours worked.
N = Number of normal hours worked
T= Total hours worked to make 165

The information provided is:

Price for THE FIRST 40 hours is 3.
Price for extra hours is 4.5
At least the total income should be 165
N=40


Then,
N*3 + 4.5*E = 165
40*3 + 4.5*E = 165
4.5*E = 165 - 40*3
4.5*E = 165 - 120
4.5*E = 45
E = 45/4.5
E = 450/45
E = 10

So, the the formula of the total hours worked to make at least 165 is:
T = N + E
T = 40+10
T = 50
Answer: To make at least 165, June has to work 50 hours.